# 2cm2cm1cm1cm

### v4

@@ 2.6 Bach, Lieb and Solovej’s Approach to Quadratic Hamiltonians
Let $\mathrm{H}$ be a quadratic Hamiltonian with respect to Definition 2.10. Then, (37) defines a strongly continuous group $(U_{t}=\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations on $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$, where $h$ is a self-adjoint operator acting on $\mathcal{H}$ satisfying $h=h^{\ast}=-\mathfrak{A}h\mathfrak{A}$ with domain $\mathcal{D}(h)=\mathfrak{A}(\mathcal{D}(h))\subseteq\mathcal{\mathcal{H}}$.
Proof. As done in [15, Equation (2b.36)], for any quadratic Hamiltonian $\mathrm{H}$ with respect to Definition 2.10, one computes from the CAR (20) and $U_{t}\mathfrak{A=A}U_{t}$ that, for any $\varphi,\psi\in\mathcal{H}$ and $t\in\mathbb{R}$,
[…]
where $\left\{A,B\right\}\doteq AB+BA$ is the usual anticommutator. So, by elementary computations using the strong continuity of the family $\left(\mathrm{e}^{it\mathrm{H}/2}\right)_{t\in\mathbb{R}}$ of unitary operators, we deduce that $(U_{t})_{t\in\mathbb{R}}$ is a strongly continuous one-parameter group of Bogoliubov transformations, see [31, Chapter I, Theorem 5.8]. In particular, we infer from Stone’s theorem [44, Theorem 6.2] the existence of some self-adjoint operator $h=h^{\ast}$ acting on $\mathcal{H}$ such that $U_{t}=\mathrm{e}^{ith}$ for any $t\in\mathbb{R}$. The assertion then follows from Lemma 3.13.
@@
Take a compatible quadratic Hamiltonian $\mathrm{H}$ with respect to Definitions 2.10 and 2.13. If $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ in (39) then, for some constant $E_{0}\in\mathbb{R}$,
@@
see Equations (169)–(170) for more details. Using the group property of $(\mathrm{e}^{it\mathrm{H}/2})_{t\in\mathbb{R}}$ and $(\mathrm{e}^{it\mathrm{H}_{0}/2})_{t\in\mathbb{R}}$ note that
$c_{t+s}\mathrm{e}^{i\left(t+s\right)\mathrm{H}_{0}/2}=\mathrm{e}^{i\left(t+s\right)\mathrm{H}/2}=c_{t}c_{s}\mathrm{e}^{i\left(t+s\right)\mathrm{H}_{0}/2}\ ,\qquad s,t\in{\mathbb{R}}\ .$ (43)
Additionally, for any $s,t\in{\mathbb{R}}$ and every normalized vector $\varphi\in\mathcal{F}$, i.e., vectors $\varphi$ in the Fock space $\mathcal{F}$ satisfying $\left\|\varphi\right\|_{\mathcal{F}}=1$,
$\left|c_{t}\right|=\left\|c_{t}\varphi\right\|_{\mathcal{F}}=\left\|\mathrm{e}^{-it\mathrm{H}_{0}/2}\mathrm{e}^{it\mathrm{H}/2}\varphi\right\|_{\mathcal{F}}=\left\|\varphi\right\|_{\mathcal{F}}=1$ (44)
as well as
$\left|c_{t}-c_{s}\right|$ $=$ $\left\|\left(c_{t}-c_{s}\right)\varphi\right\|_{\mathcal{F}}\leq\left\|\left(\mathrm{e}^{-it\mathrm{H}_{0}/2}\mathrm{e}^{it\mathrm{H}/2}-\mathrm{e}^{-is\mathrm{H}_{0}/2}\mathrm{e}^{is\mathrm{H}/2}\right)\varphi\right\|_{\mathcal{F}}$ (45)
$\leq$ $\left\|\mathrm{e}^{-it\mathrm{H}_{0}/2}\right\|_{\mathrm{op}}\left\|\left(\mathrm{e}^{it\mathrm{H}/2}-\mathrm{e}^{is\mathrm{H}/2}\right)\varphi\right\|_{\mathcal{F}}+\left\|\left(\mathrm{e}^{-it\mathrm{H}_{0}/2}-\mathrm{e}^{-is\mathrm{H}_{0}/2}\right)\mathrm{e}^{is\mathrm{H}/2}\varphi\right\|_{\mathcal{F}}\ ,$
thanks to the triangle inequality. Since $(\mathrm{e}^{it\mathrm{H}/2})_{t\in\mathbb{R}}$ and $(\mathrm{e}^{it\mathrm{H}_{0}/2})_{t\in\mathbb{R}}$ are strongly continuous groups of unitary operators, it follows from (43)–(45) that $(c_{t})_{t\in\mathbb{R}}\in C\left(\mathbb{R};\mathbb{C}\right)$ is a continuous complex-valued function satisfying
$\left|c_{t}\right|=1\qquad\text{and}\qquad c_{t+s}=c_{t}c_{s}$
for any $s,t\in{\mathbb{R}}$. As is well-known, there is therefore a real constant $\theta\in\mathbb{R}$ such that $c_{t}=\mathrm{e}^{it\theta/2}$ for any $t\in{\mathbb{R}}$. Combining this last property with (42) and the uniqueness of the generator of the strongly continuous group $\left(\mathrm{e}^{it\mathrm{H}/2}=\mathrm{e}^{it\left(\mathrm{H}_{0}+\theta\right)/2}\right)_{t\in\mathbb{R}}$ of unitary operators (see [44, Theorem 6.2]), we arrive at the assertion.
Corollary 2.14 is of course the converse of Corollary 2.11 in the Hilbert-Schmidt case. Note, however, that it is not clear whether the off-diagonal operator $D_{0}$ in Definition 2.13 (see (39)) is a bounded operator while it must be a Hilbert-Schmidt operator for the quadratic Hamiltonians defined in Section 2.2, as seen in Corollaries 2.11 and 2.14.
On the one hand, the Hilbert-Schmidt condition for $D_{0}$ results from the (natural) use of the domain $\mathcal{D}_{0}$ (5), which is a core for all our quadratic Hamiltonians. See Proposition 3.4. In particular, the vacuum state $\Psi\doteq\left(1,0,\ldots\right)$ of the fermionic Fock space $\mathcal{F}$ always belongs to the domain of our quadratic Hamiltonians, see Equation (7). By contrast, $\Psi$ may not be in the domain of a quadratic Hamiltonian with respect to Definition 2.10. But, a quadratic Hamiltonian $\mathrm{H}$ of Definition 2.10 that does not include $\Psi$ in its domain would be intrinsically different from our quadratic Hamiltonians and would refer to an abnormal model in Physics. We strongly suspect that the vacuum state would not belong to the domain of H when the operator $D_{0}$ of Definition 2.13 (see (39)) is not a Hilbert-Schmidt operator. This conjecture is based on our next theorems.
@@
Take any compatible quadratic Hamiltonian $\mathrm{H}$ with respect to Definitions 2.10 and 2.13, the domain of which is $\mathcal{D}\left(\mathrm{H}\right)\subseteq\mathcal{F}$. Assume that, for some $c\in\mathbb{R}^{+}\cup\{\infty\}$,
@@
Proof. Combine Corollaries 2.14 and 3.27 (ii), keeping in mind Definitions 2.10 and 2.13.
@@
In applications, it is very natural to have an off-diagonal operator-valued coefficient $\overline{D}_{0}$ that is relatively bounded with respect of the diagonal one $\Upsilon_{0}$. For the block operator matrix (39) with densely defined linear operators $\Upsilon_{0}$ and $\overline{D}_{0}$ acting on $\mathfrak{h}$, it means that $\mathcal{D}(\overline{D}_{0})\supseteq\mathcal{D}(\Upsilon_{0})$ and that there exist constants $\alpha_{1},\alpha_{2}\in\mathbb{R}^{+}$ such that, for any $\varphi\in\mathcal{D}\left(\Upsilon_{0}\right)$,
@@
See [45, Equation (X.19a)]. It is an usual assumption which enables well-defined block operator matrices like (39). See for instance [37, 43, 46]. In Physics, it means that the off-diagonal terms $D_{0}$,$\overline{D}_{0}$ can be viewed as a bounded perturbation of the free system, characterized  by the free (one particle) Hamiltonian $\Upsilon_{0}$. On the level of quadratic Hamiltonians, it means that the non-$\mathrm{N}$–diagonal part (see (95)) should be relatively bounded with respect to the $\mathrm{N}$–diagonal part of the given quadratic Hamiltonian. The latter is in fact true whenever $D_{0}$ is not only relatively bounded with respect $\Upsilon_{0}$ in the operator norm as in (47) but in the Hilbert-Schmidt norm, in the sense of Lemma 3.1 with $\theta=\Upsilon_{0}+\mu\mathbf{1}\geq 0$ for some appropriate $\mu\in\mathbb{R}$. The Shale-Stinespring-like condition for such generators can be studied by applying directly Theorem 2.15:
@@
Take any compatible quadratic Hamiltonian $\mathrm{H}$ with respect to Definitions 2.10 and 2.13, the domain of which is $\mathcal{D}\left(\mathrm{H}\right)\subseteq\mathcal{F}$. Assume that $D_{0}$ is relatively bounded with respect of $\Upsilon_{0}$. Then, $\Psi\doteq\left(1,0,\ldots\right)\in\mathcal{D}\left(\mathrm{H}\right)$ iff $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$.
@@
Similar to the usual Shale-Stinespring condition given by Theorem 2.8, the property $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ in Theorem 2.15 and Corollary 2.17 can be seen as a Shale-Stinespring condition for generators of implementable groups of Bogoliubov transformations. This condition is directly related with the question of whether the vacuum state belongs to the domain of the quadratic Hamiltonian or not. When it is the case, we obtain the quadratic Hamiltonians diagonalized in the present paper.
@@
Take any compatible quadratic Hamiltonian $\mathrm{H}$ with respect to Definitions 2.10 and 2.13, the domain of which is $\mathcal{D}\left(\mathrm{H}\right)\subseteq\mathcal{F}$. Assume $D_{0}\in\mathcal{B}(\mathfrak{h})$. Then, the following assertions hold true:
@@
    $(\varkappa_{t}(D_{0}))_{t\in\mathbb{R}}\subseteq\mathcal{B}(\mathfrak{h})$ belongs to $\mathcal{L}^{2}(\mathfrak{h})$ for times $t$ in a neighborhood of 00 and is even continuous in $\mathcal{L}^{2}(\mathfrak{h})$ at $t=0$.
@@
3 Technical Results
Before starting the proofs, we add a few general definitions. First, $s\wedge t$ and $s\vee t$ stand respectively for the minimum and the maximum of real numbers $s,t\in\mathbb{R}$. This notation has not to be mixed with the exterior product defining the subspace $\wedge^{n}\mathfrak{h}$ of totally antisymmetric $n$–particle wave functions in $\mathfrak{h}^{\otimes n}$.
The Banach space of bounded operators from a Banach space $\left(\mathcal{X},\left\|\cdot\right\|_{\mathcal{X}}\right)$ to another Banach space $\left(\mathcal{Y},\left\|\cdot\right\|_{\mathcal{Y}}\right)$ is denoted by $\mathcal{B}(\mathcal{X},\mathcal{Y})$, with operator norm
@@ 3.1 Self-adjointness of Quadratic Operators
$\left\|\mathbf{D}\varphi\right\|_{\mathcal{F}}^{2}\leq\left\|\theta^{-1/2}D_{0}\theta^{-1/2}\right\|_{2}^{2}\sum_{k,l\in\mathbb{N}}\theta_{k}\theta_{l}\left\langle a_{l}\varphi,a_{k}^{\ast}a_{k}a_{l}\varphi\right\rangle_{\mathcal{F}}\\ =\left\|\theta^{-1/2}D_{0}\theta^{-1/2}\right\|_{2}^{2}\sum_{k,l\in\mathbb{N}}\left(1-\delta_{k,l}\right)\left\langle\theta_{l}a_{l}^{\ast}a_{l}\varphi,\theta_{k}a_{k}^{\ast}a_{k}\varphi\right\rangle_{\mathcal{F}}\leq\left\|\theta^{-1/2}D_{0}\theta^{-1/2}\right\|_{2}^{2}\left\|d\Gamma\left(\theta\right)\varphi\right\|_{\mathcal{F}}^{2}\ ,$
@@
where $\mathcal{D}\left(\Upsilon_{0}\right)\subseteq\mathfrak{h}$ is the (dense) domain of $\Upsilon_{0}=\Upsilon_{0}^{\ast}$, $\wedge^{0}\mathcal{D}\left(\Upsilon_{0}\right)\doteq\mathbb{C}$ while for $n\in\mathbb{N}$, $\wedge^{n}\mathcal{D}\left(\Upsilon_{0}\right)$ is the (dense) subspace of totally antisymmetric $n$–particle wave functions in $\mathcal{D}\left(\Upsilon_{0}\right)^{\otimes n}$, the $n$–fold tensor product of $\mathcal{D}\left(\Upsilon_{0}\right)^{\otimes n}\subseteq\mathfrak{h}^{\otimes n}$.
@@
[…]
$\varphi_{\pm\epsilon}\doteq\left(1\pm i\epsilon d\Gamma(\Upsilon_{0})\right)^{-1}\varphi\ ,$
[…]
$z\left\|P_{n}\varphi\right\|_{\mathcal{F}}^{2}=z\langle P_{n}\varphi,\varphi\rangle_{\mathcal{F}}=\lim_{\epsilon\rightarrow 0}z\langle P_{n}\varphi_{\epsilon},\varphi\rangle_{\mathcal{F}}=\lim_{\epsilon\rightarrow 0}\langle P_{n}\varphi_{\epsilon},\mathrm{H}_{0}^{\ast}\varphi\rangle_{\mathcal{F}}=\lim_{\epsilon\rightarrow 0}\langle\mathrm{H}_{0}P_{n}\varphi_{\epsilon},\varphi\rangle_{\mathcal{F}}$
and, in the same way,
$\overline{z}\left\|P_{n}\varphi\right\|_{\mathcal{F}}^{2}=\lim_{\epsilon\rightarrow 0}\langle\varphi,\mathrm{H}_{0}P_{n}\varphi_{-\epsilon}\rangle_{\mathcal{F}}\ .$
So, by summing and using (49)–(51) we obtain that
$2i\mathrm{Im}\left(z\left\|P_{n}\varphi\right\|_{\mathcal{F}}^{2}\right)$ $=$ $\lim_{\epsilon\rightarrow 0}\left\{\langle d\Gamma(\Upsilon_{0})P_{n}\varphi_{\epsilon},\varphi\rangle_{\mathcal{F}}-\langle\varphi,d\Gamma(\Upsilon_{0})P_{n}\varphi_{-\epsilon}\rangle_{\mathcal{F}}\right.$
$\left.+\langle(\mathbf{D}+\mathbf{D}^{\ast})P_{n}\varphi_{\epsilon},\varphi\rangle_{\mathcal{F}}-\langle\varphi,(\mathbf{D}+\mathbf{D}^{\ast})P_{n}\varphi_{-\epsilon}\rangle_{\mathcal{F}}\right\}$
[…]
$2i\mathrm{Im}\left(z\left\|P_{n}\varphi\right\|_{\mathcal{F}}^{2}\right)$ $=$ $\langle(\mathbf{D}+\mathbf{D}^{\ast})P_{n}\varphi,\varphi\rangle_{\mathcal{F}}-\langle\varphi,(\mathbf{D}+\mathbf{D}^{\ast})P_{n}\varphi\rangle_{\mathcal{F}}$
$=$ $\langle\mathbf{D}P_{n}\varphi,P_{n-2}\varphi\rangle_{\mathcal{F}}+\langle\mathbf{D}^{\ast}P_{n}\varphi,P_{n+2}\varphi\rangle_{\mathcal{F}}$
$-\langle\mathbf{D}P_{n+2}\varphi,P_{n}\varphi\rangle_{\mathcal{F}}-\langle\mathbf{D}^{\ast}P_{n-2}\varphi,P_{n}\varphi\rangle_{\mathcal{F}}$
[…]
$2i\mathrm{Im}\left(z\sum_{n=0}^{N}\left\|P_{n}\varphi\right\|_{\mathcal{F}}^{2}\right)$ $=$ $-\langle\mathbf{D}P_{N+2}\varphi,P_{N}\varphi\rangle_{\mathcal{F}}-\langle\mathbf{D}P_{N+1}\varphi,P_{N-1}\varphi\rangle_{\mathcal{F}}$
$+\langle\mathbf{D}^{\ast}P_{N}\varphi,P_{N+2}\varphi\rangle_{\mathcal{F}}+\langle\mathbf{D}^{\ast}P_{N-1}\varphi,P_{N+1}\varphi\rangle_{\mathcal{F}}\ .$
Then, we use the Cauchy-Schwarz inequality and apply Corollary 3.2 to arrive at the inequalities
$2\mathrm{Im}\left(z\sum_{n=0}^{N}\left\|P_{n}\varphi\right\|_{\mathcal{F}}^{2}\right)$ $\leq\left\|D_{0}\right\|_{2}\left(2(N+2)\left\|P_{N}\varphi\right\|_{\mathcal{F}}\left\|P_{N+2}\varphi\right\|_{\mathcal{F}}+2(N+1)\left\|P_{N-1}\varphi\right\|_{\mathcal{F}}\left\|P_{N+1}\varphi\right\|_{\mathcal{F}}\right)$
$\leq(N+2)\left\|D_{0}\right\|_{2}\left(\left\|P_{N-1}\varphi\right\|_{\mathcal{F}}^{2}+\left\|P_{N}\varphi\right\|_{\mathcal{F}}^{2}+\left\|P_{N+1}\varphi\right\|_{\mathcal{F}}^{2}+\left\|P_{N+2}\varphi\right\|_{\mathcal{F}}^{2}\right)\ .$ (54)
(Note that the last upper bound is obtained from the trivial inequality $2xy\leq x^{2}+y^{2}$ for $x,y\in\mathbb{R}_{0}^{+}$.) By contradiction, suppose that $\varphi\neq 0$. Necessarily there exist $N_{0}\in\mathbb{N}$ and a constant $C\in\mathbb{R}^{+}$ such that for all natural numbers $N\geq N_{0}$,
$\sum_{n=0}^{N}\left\|P_{n}\varphi\right\|_{\mathcal{F}}^{2}\geq\sum_{n=0}^{N_{0}}\left\|P_{n}\varphi\right\|_{\mathcal{F}}^{2}=C>0\ .$
As a result, for all natural numbers $N\geq N_{0}$, we deduce from (54) that
$\frac{2C\mathrm{Im}z}{N+2}\leq\left\|D_{0}\right\|_{2}\sum_{j=N-1}^{N+2}\left\|P_{j}\varphi\right\|_{\mathcal{F}}^{2}\ .$
If we sum up to $M\in\mathbb{N}$ with $M\geq N_{0}$ we obtain
$\sum_{N=N_{0}}^{M}\frac{2C\mathrm{Im}z}{N+2}\leq\left\|D_{0}\right\|_{2}\sum_{N=N_{0}}^{M}\sum_{j=N-1}^{N+2}\left\|P_{j}\varphi\right\|_{\mathcal{F}}^{2}\ .$
[…]

### v6

@@ 2.6 Bach, Lieb and Solovej’s Approach to Quadratic Hamiltonians
Let $\mathrm{H}$ be a quadratic Hamiltonian in the sense of Definition 2.10. Then, (37) defines a strongly continuous group $(\mathcal{U}_{t}=\mathrm{e}^{ith})_{t\in\mathbb{R}}$ of Bogoliubov transformations on $\mathcal{H}\doteq\mathfrak{h}\oplus\mathfrak{h}$, where $h$ is a self-adjoint operator acting on $\mathcal{H}$ satisfying $h=h^{\ast}=-\mathfrak{A}h\mathfrak{A}$ with domain $\mathcal{D}(h)=\mathfrak{A}(\mathcal{D}(h))\subseteq\mathcal{\mathcal{H}}$.
Proof. As done in [15, Equation (2b.36)], for any quadratic Hamiltonian $\mathrm{H}$ in the sense of Definition 2.10, one computes from the CAR (20) and $\mathcal{U}_{t}\mathfrak{A=A}\mathcal{U}_{t}$ that, for any $\varphi,\psi\in\mathcal{H}$ and $t\in\mathbb{R}$,
$\left\{\mathbf{\pi}_{\mathcal{F}}\left(\mathrm{B}\left(\varphi\right)\right),\mathrm{e}^{it\mathrm{H}/2}\mathbf{\pi}_{\mathcal{F}}\left(\mathrm{B}\left(\mathfrak{A}\psi\right)\right)\mathrm{e}^{-it\mathrm{H}/2}\right\}=\mathbf{\pi}_{\mathcal{F}}\left(\left\{\mathrm{B}\left(\varphi\right),\mathrm{B}\left(\mathcal{U}_{t}\mathfrak{A}\psi\right)\right\}\right)=\left\langle\varphi,\mathcal{U}_{t}\psi\right\rangle_{\mathcal{H}}\,\mathbf{1}\ ,$
where $\left\{A,B\right\}\doteq AB+BA$ is the usual anticommutator. So, by elementary computations using the strong continuity of the family $\left(\mathrm{e}^{it\mathrm{H}/2}\right)_{t\in\mathbb{R}}$ of unitary operators, we deduce that $(\mathcal{U}_{t})_{t\in\mathbb{R}}$ is a strongly continuous one-parameter group of Bogoliubov transformations, see [31, Chapter I, Theorem 5.8]. In particular, we infer from Stone’s theorem [44, Theorem 6.2] the existence of some self-adjoint operator $h=h^{\ast}$ acting on $\mathcal{H}$ such that $\mathcal{U}_{t}=\mathrm{e}^{ith}$ for any $t\in\mathbb{R}$. The assertion then follows from Lemma 4.1.
@@
Take a compatible quadratic Hamiltonian $\mathrm{H}$ in the sense of Definitions 2.10 and 2.13. If $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ in (39) then, for some constant $E_{0}\in\mathbb{R}$,
@@
see Equations (165)–(166) for more details. Since $(\mathrm{e}^{it\mathrm{H}/2})_{t\in\mathbb{R}}$ and $(\mathrm{e}^{it\mathrm{H}_{0}/2})_{t\in\mathbb{R}}$ are strongly continuous groups of unitary operators, using (42) and elementary estimates, one shows [^15] For more details, see [45]. that $(c_{t})_{t\in\mathbb{R}}\in C\left(\mathbb{R};\mathbb{C}\right)$ is a continuous complex-valued function satisfying $\left|c_{t}\right|=1$ and $c_{t+s}=c_{t}c_{s}$ for any $s,t\in{\mathbb{R}}$. As is well-known, there is therefore a real constant $\theta\in\mathbb{R}$ such that $c_{t}=\mathrm{e}^{it\theta/2}$ for any $t\in{\mathbb{R}}$. Combining this last property with (42) and the uniqueness of the generator of the strongly continuous group $\left(\mathrm{e}^{it\mathrm{H}/2}=\mathrm{e}^{it\left(\mathrm{H}_{0}+\theta\right)/2}\right)_{t\in\mathbb{R}}$ of unitary operators (see [44, Theorem 6.2]), we arrive at the assertion.
Corollary 2.14 is of course the converse of Corollary 2.11 in the Hilbert-Schmidt case. Note, however, that it is not clear whether the off-diagonal operator $D_{0}$ in Definition 2.13 (see (39)) is a bounded operator.
On the one hand, the Hilbert-Schmidt condition for $D_{0}$ results from the (natural) use of the domain $\mathcal{D}_{0}$ (5), which is a core for all our quadratic Hamiltonians. See Proposition 3.4. In particular, the vacuum state $\Psi\doteq\left(1,0,\ldots\right)$ of the fermionic Fock space $\mathcal{F}$ always belongs to the domain of our quadratic Hamiltonians, see Equation (7). By contrast, $\Psi$ may not be in the domain of a quadratic Hamiltonian in the sense of Definition 2.10. But, a quadratic Hamiltonian $\mathrm{H}$ of Definition 2.10 that does not include $\Psi$ in its domain would be intrinsically different from our quadratic Hamiltonians and would refer to an non-traditional (albeit possibly very interesting) model in Physics. We strongly suspect that the vacuum state would not belong to the domain of H when the operator $D_{0}$ of Definition 2.13 (see (39)) is not a Hilbert-Schmidt operator. This conjecture is based on our next theorems.
@@
Take any compatible quadratic Hamiltonian $\mathrm{H}$ in the sense of Definitions 2.10 and 2.13, the domain of which is $\mathcal{D}\left(\mathrm{H}\right)\subseteq\mathcal{F}$. Assume that, for some $c\in\mathbb{R}^{+}\cup\{\infty\}$,
@@
Proof. Combine Corollaries 2.14 and 4.15 (ii), keeping in mind Definitions 2.10 and 2.13.
@@
In applications, it is very natural to have an off-diagonal operator-valued coefficient $\overline{D}_{0}$ that is relatively bounded with respect to the diagonal one $\Upsilon_{0}$. For the block operator matrix (39) with densely defined linear operators $\Upsilon_{0}$ and $\overline{D}_{0}$ acting on $\mathfrak{h}$, it means that $\mathcal{D}(\overline{D}_{0})\supseteq\mathcal{D}(\Upsilon_{0})$ and that there exist constants $\alpha_{1},\alpha_{2}\in\mathbb{R}^{+}$ such that, for any $\varphi\in\mathcal{D}\left(\Upsilon_{0}\right)$,
@@
See [46, Equation (X.19a)]. It is an usual assumption which enables well-defined block operator matrices like (39). See for instance [37, 43, 47]. In Physics, it means that the off-diagonal terms $D_{0}$$\overline{D}_{0}$ can be viewed as a bounded perturbation of the free system, characterized by the free (one particle) Hamiltonian $\Upsilon_{0}$. On the level of quadratic Hamiltonians, it means that the non-$\mathrm{N}$–diagonal part (see (91)) should be relatively bounded with respect to the $\mathrm{N}$–diagonal part of the given quadratic Hamiltonian. The latter is in fact true whenever $D_{0}$ is not only relatively bounded with respect $\Upsilon_{0}$ in the operator norm as in (44) but in the Hilbert-Schmidt norm, in the sense of Lemma 3.1 with $\theta=\Upsilon_{0}+\mu\mathbf{1}\geq 0$ for some appropriate $\mu\in\mathbb{R}$. The Shale-Stinespring-like condition for such generators can be studied by applying directly Theorem 2.15:
@@
Take any compatible quadratic Hamiltonian $\mathrm{H}$ in the sense of Definitions 2.10 and 2.13, the domain of which is $\mathcal{D}\left(\mathrm{H}\right)\subseteq\mathcal{F}$. Assume that $D_{0}$ is relatively bounded with respect to $\Upsilon_{0}$. Then, $\Psi\doteq\left(1,0,\ldots\right)\in\mathcal{D}\left(\mathrm{H}\right)$ iff $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$.
@@
Similar to the usual Shale-Stinespring condition given by Theorem 2.8, the property $D_{0}\in\mathcal{L}^{2}(\mathfrak{h})$ in Theorem 2.15 and Corollary 2.17 can be seen as a Shale-Stinespring condition [^17] The Shale-Stinespring condition on Bogoliubov transformations does not necessarily imply a Hilbert-Schmidt condition on the off-diagonal elements of its generator. See, however, Theorem 2.18 (i). for generators of implementable groups of Bogoliubov transformations. This condition is directly related with the question of whether the vacuum state belongs to the domain of the quadratic Hamiltonian or not. When it is the case, we obtain the quadratic Hamiltonians diagonalized in the present paper.
@@
Take any compatible quadratic Hamiltonian $\mathrm{H}$ in the sense of Definitions 2.10 and 2.13, the domain of which is $\mathcal{D}\left(\mathrm{H}\right)\subseteq\mathcal{F}$. Assume $D_{0}\in\mathcal{B}(\mathfrak{h})$. Then, the following assertions hold true:
@@
    $(\varkappa_{t}(D_{0}))_{t\in\mathbb{R}}\subseteq\mathcal{B}(\mathfrak{h})$ belongs to $\mathcal{L}^{2}(\mathfrak{h})$ for times $t$ in a neighborhood of 0 and is even continuous in $\mathcal{L}^{2}(\mathfrak{h})$ at $t=0$.
@@
3 Technical Results: Implementation of the Brockett-Wegner Flow
The technical results are divided into two parts. In this part, after proving some preliminary results, we study the implementation of the Brockett-Wegner flow (Section 2.5) in order to ultimately prove Theorem 2.4. Before starting the proofs, we add a few general definitions that are used in both technical sections.
First, $s\wedge t$ and $s\vee t$ stand respectively for the minimum and the maximum of real numbers $s,t\in\mathbb{R}$. This notation has not to be mixed with the exterior product defining the subspace $\wedge^{n}\mathfrak{h}$ of totally antisymmetric nn–particle wave functions in $\mathfrak{h}^{\otimes n}$. The Banach space of bounded operators from a Banach space $\left(\mathcal{X},\left\|\cdot\right\|_{\mathcal{X}}\right)$ to another Banach space $\left(\mathcal{Y},\left\|\cdot\right\|_{\mathcal{Y}}\right)$ is denoted by $\mathcal{B}(\mathcal{X},\mathcal{Y})$, with operator norm
@@ 3.1 Self-adjointness of Quadratic Operators
$\left\|\mathbf{D}\varphi\right\|_{\mathcal{F}}^{2}\leq\left\|\theta^{-1/2}D_{0}\theta^{-1/2}\right\|_{2}^{2}\sum_{k,l\in\mathbb{N}}\theta_{k}\theta_{l}\left\langle a_{l}\varphi,a_{k}^{\ast}a_{k}a_{l}\varphi\right\rangle_{\mathcal{F}}\\ =\left\|\theta^{-1/2}D_{0}\theta^{-1/2}\right\|_{2}^{2}\sum_{k,l\in\mathbb{N}}\left(1-\delta_{k,l}\right)\left\langle\theta_{l}a_{l}^{\ast}a_{l}\varphi,\theta_{k}a_{k}^{\ast}a_{k}\varphi\right\rangle_{\mathcal{F}}\leq\left\|\theta^{-1/2}D_{0}\theta^{-1/2}\right\|_{2}^{2}\left\|d\Gamma\left(\theta\right)\varphi\right\|_{\mathcal{F}}^{2}\ ,\$
@@
where $\mathcal{D}\left(\Upsilon_{0}\right)\subseteq\mathfrak{h}$ is the (dense) domain of $\Upsilon_{0}=\Upsilon_{0}^{\ast}$, $\wedge^{0}\mathcal{D}\left(\Upsilon_{0}\right)\doteq\mathbb{C}$ while for $n\in\mathbb{N}$, $\wedge^{n}\mathcal{D}\left(\Upsilon_{0}\right)$ is the (dense) subspace of totally antisymmetric nn–particle wave functions in $\mathcal{D}\left(\Upsilon_{0}\right)^{\otimes n}$, the nn–fold tensor product of $\mathcal{D}\left(\Upsilon_{0}\right)^{\otimes n}\subseteq\mathfrak{h}^{\otimes n}$.
@@
Proof. The proof is the same as the one of [7, Theorem 5.3] (itself inspired by [11, Theorem 6.1]), up to obvious modifications, and is thus omitted. In particular, one shows $\ker\left(\mathrm{H}_{0}^{\ast}\pm i\right)=\left\{0\right\}$, which is a well-known criterion for essential self-adjointness (see, e.g., [25, Theorem VIII.3 and its corollary] or [44, Proposition 3.8]). See [45] for the detailed proof in this case.